Bond Yields [LO2] Seether Co. wants to issue new 20-year bonds needed expansion projects. The company currently has 8 percent coupon bonds on the market that sell for $930, make semiannual payments, and mature in 20 years. What coupon rate should the company set on its new bonds if it wants them to sell at par? 19.
Answer #1

Answer: Company can set coupon rate on its new bonds is same as required rate of return. Required rate of return is nothing but the YTM.

YTM = [C + (F-P) / n] / (F+P)/2

C : 1000*8%*1/2 = $40, F = $1000, P = $930, Mode = Semiannual, n: 20*2 = 40 periods

Putting all the above values in the formula, we get:

YTM = [40 +(1000-930) / 40] / (1000+930) / 2

YTM = 4.37%

As the coupon payment is semiannual so this rate will be multiplied by 2 so the YTM = 4.37 * 2

YTM = 8.74% is the final answer.